steve816:

Determine the convergence or divergence of the series. This one I have no clue what to do.

steve816:

\[\sum_{k=1}^{\infty}\frac{ k \ln k }{ (k+1)^3 }\]

satellite73:

you need a test?

satellite73:

compare it to \(\sum\frac{ln(k)}{k^2}\) which converges by the integral test

steve816:

yay, that's smart.

steve816:

I figured it out with that comparison.